# Generating bilingual example sentences with large language models as lexicography assistants

### v1

@@ Abstract
We present a study of LLMs’ performance in generating and rating example sentences for bilingual dictionaries across languages with varying resource levels: French (high-resource), Indonesian (mid-resource), and Tetun (low-resource), with English as the target language. We evaluate the quality of LLM-generated examples against the GDEX (Good Dictionary EXample) criteria: typicality, informativeness, and intelligibility (Kilgarriff et al., 2008). Our findings reveal that while LLMs can generate reasonably good dictionary examples, their performance degrades significantly for lower-resourced languages. We also observe high variability in human preferences for example quality, reflected in low inter-annotator agreement rates. To address this, we demonstrate that in-context learning can successfully align LLMs with individual annotator preferences. Additionally, we explore the use of pre-trained language models for automated rating of examples, finding that sentence perplexity serves as a good proxy for typicality and intelligibility in higher-resourced languages. Our study also contributes a novel dataset of 600 ratings for LLM-generated sentence pairs, and provides insights into the potential of LLMs in reducing the cost of lexicographic work, particularly for low-resource languages.
@@ 1 Introduction
Example sentences in bilingual dictionaries play a crucial role in language learning, helping both to understand the meaning of headwords (words that mark a separate entry in the dictionary), and their usage in context Potgieter (2012); Nielsen (2014); Caballero (2024). What makes candidate sentences good as examples is the subject of linguistic research, with Kilgarriff et al. (2008) proposing the GDEX (Good Dictionary EXample) framework, which qualifies good examples as typical ("exhibiting frequent and well-dispersed patterns of usage"), intelligible ("avoiding gratuitously difficult lexis and structures"), and informative ("helping to elucidate the definition"), as illustrated in Table 1.
@@
  • A novel dataset of 600 sentence ratings for LLM-generated example sentence pairs in French, Indonesian, and Tetun as source, and English as target. Each pair is rated against 5 criteria, resulting in 3,000 individual annotations.
@@ 2 Background
Table 2: Example LLM-generated sentences their associated ratings from annotators for languages covered in this study.
@@
Despite the growing body of research on LLMs in lexicography, several areas remain unexplored. First, there has been no structured evaluation of LLM capabilities in generating example sentences for bilingual dictionaries. Second, the potential of LLMs to help assess the quality of examples in a bilingual context has not been systematically investigated. Lastly, we have not found comprehensive studies examining LLM-based optimisation techniques—such as prompt engineering, fine-tuning, and in-context learning—for the specific task of generating dictionary examples. Addressing these research gaps could advance our understanding of how to effectively harness LLMs for creating high-quality, contextually appropriate example sentences in bilingual dictionaries, across languages of varying resource levels.
@@ 3.1 Methodology for generation
fra 4.79 4.57 3.06*
ind 4.36 4.46 -1.04
tdt 3.86 3.61 1.55
Table 3: Average overall rating for LLM-generated examples per language, with paired t-test results, where * represents a statistically significant difference between models ($p<0.05$). For rating per criteria, see the distribution bar plot in Figure 2.
@@
Figure 2: Rating distributions for GDEX criteria and translation correctness for French, Indonesian, and Tetun
@@ 3.2 Quality of LLM-generated examples
Comparing overall rating for the two LLMs used in the study, we find that GPT-4o outperforms Llama3.1 for French (4.79 vs. 4.57), with a statistically significant t-statistic of over 3 indicating a substantial difference between the two models relative to variation in the data. For Indonesian and Tetun however, the paired t-test indicated that the difference between the two models is not statistically significant compared to the variation in the data. We therefore observe variability in LLM output quality that is uneven across language depending on resource level and shows that performance degradation is not always predictable from resource level.
@@
fra $4.74$ $4.62$ $1.830$
ind $4.09$ $4.73$ $-6.273$*
tdt $3.62$ $3.85$ $-1.909$
Table 4: Average rating per annotator with paired t-test results, where * represents a statistically significant difference between annotators ($p<0.05$). For each language, A1 is the annotator with a computational linguistics background.
@@ 3.3 A note on inter-annotator agreement
We note that low inter-annotator agreement for rating examples was observed in previous studies Benedetti et al. (2024b). This finding guides our further experiments: (1) when working with in-context learning, we favour aligning LLM rating with one annotator’s judgement at a time, rather than aligning with contradicting ratings coming from multiple annotators (Section 4.1); (2) when working with pretrained language models, which are not fine-tuned to annotator preference, we only measure alignment with the annotator who has a computational linguistics background (Section 4.2).
@@ 4 Automated rating of example sentences
This section describes our experiments for automating the rating of LLM-generated sentences, in alignment with human ratings.
@@ 4.2 Rating through pre-trained language models
To test the hypotheses, we use pre-trained encoder-only language models: CamemBERT-large for French Martin et al. (2019), IndoBERT for Indonesian Koto et al. (2020). For Tetun, given the absence of existing encoder-only models for the language, we fine-tune XLM-RoBERTa-large Conneau et al. (2019) on MADLAD-400 Kudugunta et al. (2023) which is the largest Tetun monolingual corpus available, using the hyperparameters in Adelani et al. (2021). We release the weights of this model for future researchers.
@@ Acknowledgments
We would like to extend our gratitude to Professor Hanna Suominen for her valuable feedback and guidance throughout this study. We also thank Gabriel de Jesus and his team, Isabel Pereira (Catalpa International), Tungga Dewi, and Matilda Merx for their support in data collection and annotation. This research was supported by The University of Melbourne’s Research Computing Services and the Petascale Campus Initiative.

### v2

@@ Abstract
We present a study of LLMs’ performance in generating and rating example sentences for bilingual dictionaries across languages with varying resource levels: French (high-resource), Indonesian (mid-resource), and Tetun (low-resource), with English as the target language. We evaluate the quality of LLM-generated examples against the GDEX (Good Dictionary EXample) criteria: typicality, informativeness, and intelligibility (Kilgarriff et al., 2008). Our findings reveal that while LLMs can generate reasonably good dictionary examples, their performance degrades significantly for lower-resourced languages. We also observe high variability in human preferences for example quality, reflected in low inter-annotator agreement rates. To address this, we demonstrate that in-context learning can successfully align LLMs with individual annotator preferences. Additionally, we explore the use of pre-trained language models for automated rating of examples, finding that sentence perplexity serves as a good proxy for "typicality" and "intelligibility" in higher-resourced languages. Our study also contributes a novel dataset of 600 ratings for LLM-generated sentence pairs, and provides insights into the potential of LLMs in reducing the cost of lexicographic work, particularly for low-resource languages.
@@ 1 Introduction
Example sentences in bilingual dictionaries play a crucial role in language learning, helping L2 speakers to understand the meaning of headwords (words that mark a separate entry in the dictionary), and their usage in context Potgieter (2012); Nielsen (2014); Caballero (2024). What makes candidate sentences good as examples is the subject of linguistic research, with Kilgarriff et al. (2008) proposing the GDEX (Good Dictionary EXample) framework, which qualifies good examples as typical ("exhibiting frequent and well-dispersed patterns of usage"), intelligible ("avoiding gratuitously difficult lexis and structures"), and informative ("helping to elucidate the definition"), as illustrated in Table 1. In bilingual setups, the accuracy of translation between source and target examples also contributes to example quality.
@@
  • A novel dataset of 600 sentence ratings for LLM-generated example sentence pairs in French, Indonesian, and Tetun as source, and English as target. Each pair is rated against 5 criteria, resulting in 3,000 individual annotations. [^1] https://github.com/raphaelmerx/llm-bilingual-examples
@@ 2 Background
Table 2: Example LLM-generated sentences and annotator ratings for languages covered in this study.
@@
Despite the growing body of research on LLMs in lexicography, several areas remain unexplored. First, there has been no structured evaluation of LLM capabilities in generating example sentences for bilingual dictionaries, where additional challenges arise compared to monolingual dictionaries, such as maintaining GDEX criteria across languages while ensuring translation accuracy. Second, the potential of LLMs to help assess the quality of examples in a bilingual context, which could assist with example selection and with the setup of self-improvement pipelines for generation, has not been systematically investigated. Lastly, we have not found comprehensive studies examining LLM-based optimisation techniques—such as prompt engineering, fine-tuning, and in-context learning—for the specific task of generating dictionary examples. Addressing these research gaps could advance our understanding of how to effectively harness LLMs for creating high-quality, contextually appropriate example sentences in bilingual dictionaries, across languages of varying resource levels.
@@ 3.1 Methodology for generation
fra 4.79 $\pm$ 0.47 4.57 $\pm$ 0.62 $3.06$*
ind 4.36 $\pm$ 0.82 4.46 $\pm$ 0.79 $-1.04$
tdt 3.86 $\pm$ 1.18 3.61 $\pm$ 1.22 $1.55$
Table 3: Average overall rating ($\pm$ standard deviation) for LLM-generated examples per language, with paired t-test results, where * represents a statistically significant difference between models ($p<0.05$). For rating per criteria, see the distribution bar plot in Figure 2.
@@
Figure 2: Rating distributions (GPT-4o and Llama 3.1 combined) for GDEX criteria and translation correctness.
@@ 3.2 Quality of LLM-generated examples
Comparing overall rating for the two LLMs used in the study, we find that GPT-4o outperforms Llama3.1 for French (4.79 vs. 4.57), with a statistically significant t-statistic of over 3 indicating a substantial difference between the two models relative to variation in the data. For Indonesian and Tetun however, the paired t-test indicated that the difference between the two models is not statistically significant compared to the variation in the data. We therefore observe variability in LLM output quality that is uneven across languages depending on resource level and shows that performance degradation is not always predictable from resource level.
@@
fra 4.74 $\pm$ 0.56 4.62 $\pm$ 0.56 $1.830$
ind 4.09 $\pm$ 0.85 4.73 $\pm$ 0.62 $-6.273$*
tdt 3.62 $\pm$ 1.47 3.85 $\pm$ 0.88 $-1.909$
Table 4: Average rating ($\pm$ standard deviation) per annotator with paired t-test results, where * represents a statistically significant difference between annotators ($p<0.05$). For each language, A1 is the annotator with a computational linguistics background.
@@ 3.3 A note on inter-annotator agreement
We note that low inter-annotator agreement for rating examples was observed in previous studies Benedetti et al. (2024b). This finding guides our further experiments: (1) when working with in-context learning, we favour aligning LLM rating with one annotator’s judgement at a time, rather than aligning with contradicting ratings coming from multiple annotators (Section 4.1); (2) when working with pre-trained language models, which are not fine-tuned to annotator preference, we only measure alignment with the annotator who has a computational linguistics background (Section 4.2).
@@ 4 Automated rating of example sentences
Beyond baseline performance across different resource levels, we evaluate how well LLMs can assess example quality. This could enable more efficient dictionary creation pipelines, where automated rating systems that align with human judgement could help filter and select the best examples from larger sets of generated candidates, reducing the need for extensive manual review. Furthermore, reliable automated evaluation metrics could facilitate the development of self-improvement systems where LLMs learn from their own assessments to generate increasingly better examples.
@@ 4.2 Rating through pre-trained language models
To test the hypotheses, we use pre-trained encoder-only language models: CamemBERT-large for French Martin et al. (2019), IndoBERT for Indonesian Koto et al. (2020). For Tetun, given the absence of existing encoder-only models for the language, we fine-tune XLM-RoBERTa-large Conneau et al. (2019) on MADLAD-400 Kudugunta et al. (2023) which is the largest Tetun monolingual corpus available, using the hyperparameters in Adelani et al. (2021). We release the weights of this model for future researchers. [^7] https://huggingface.co/raphaelmerx/xlm-roberta-large-tetun
@@ Acknowledgments
We would like to extend our gratitude to Professor Hanna Suominen for her valuable feedback and guidance throughout this study. We also thank Gabriel de Jesus and his team, Isabel Pereira (Catalpa International), Tungga Dewi, and Matilda Merx for their support in data collection and annotation. We also thank the anonymous reviewers for their feedback. This research was supported by The University of Melbourne’s Research Computing Services and the Petascale Campus Initiative.